if the circles x2+y2-6x-8y-24=0 and 3x2+3y2+2gx+2fy+c=0 are concentric then (g,f) =
 

Grade:12

1 Answers

Arun
5 years ago
First convert them in a form.
 
x² + y² -6x - 8y -24 = 0
&
x² + y² + 2g/3 x + 2f/3 y + c/3 = 0
 
Now
 
2g/3 = -6
g = -9
 
2f/3 = -8
f = -12